OpenStudy (anonymous):

25xy^4 + 60xy^3 divided by 5xy =???? ;)

OpenStudy (anonymous):

Write it as a fraction, with 5xy in the denominator. Then try to factor out 5xy from the two terms in the numerator (you hope this works...). If you can, then you can cancel out 5xy from numerator and denominator to get the simplified expression. Does that make sense?

OpenStudy (anonymous):

kinda i really struggle with math because i never really learned the basics to math but would it be 5xy plus 12y?

OpenStudy (anonymous):

Don't give up :) You're close.. When you factor the top, what did you get?

OpenStudy (anonymous):

I got 5xy (5y^3 +12y^2)

OpenStudy (anonymous):

you "pull" the 5xy out of each of the terms in the numerator. (Pull just means dividing it out in the process of factoring).

OpenStudy (anonymous):

how did you get that because i think my process was obviously wrong but i just dont know what process to use to solve

OpenStudy (anonymous):

ok, no worries, let me build the equation... will take a minute.

OpenStudy (anonymous):

awesome thanks for taking the time to help me im blonde so that probably doesnt help me either lol

OpenStudy (anonymous):

\[\frac{ 25xy ^{4} +60xy ^{3}}{ 5xy } = \frac{ 5xy(5y ^{3}+12y ^{2} )}{ 5xy } = 5y^{3}+12y ^{2}\]

OpenStudy (anonymous):

I'll make you a deal... I'll help, you try not to think that "I can't do math". It may not be easy but it isn't hard because you're blond :)

OpenStudy (anonymous):

Anyway, how's that equation look? Can you follow the factoring in the numerator?

OpenStudy (anonymous):

When you factor something easy, like 2x + 5x, you "pull out" the common "thing" (easy non-scary math words!!). Like the common things is x, so it's x(2 + 5).

OpenStudy (anonymous):

awesome but i just dont understand what happened to the 25 and 60

OpenStudy (anonymous):

In this case, they give you a hint, sort of... the process sometimes is, look at the denominator, laugh at yourself cuz you hate fractions and think, "how can I make that fraction "go away"? Well, if it was 4/2, you would realize that 4 = 2x2, so you would cancel out the 2 on the bottom with one of the 2s on top.

OpenStudy (anonymous):

oh ok that makes sense

OpenStudy (anonymous):

Oh, well, 25 = 5 x 5 and 60 = 12 x 5 so the common number is 5... factor that out (if it was just easy algebra, like 25x + 60x, you could factor it to 5x(5 + 12) In the actual problem, x and y are also common, so factor both out so overall, you factor out 5xy on top

OpenStudy (anonymous):

It's stuff you sorta know from easier algebra, but having two variables to factor might throw you. Also, notice that actually, you could factor out y^3 from both terms on top... the only reason I didn't is that I was "clued in" by the 5xy on the bottom... that's sometimes your cue, at least on homework & tests... try to factor that out.

OpenStudy (anonymous):

They could give you harder looking problems too that aren't really harder... you might have to factor the bottom first like if they gave you 10xy+5xy^2, you could factor to get 5xy(1 + y^2)... and then try to factor either one of those terms out of the numerator to allow you to cancel terms in the fraction (boo, fractions, remember!)

OpenStudy (anonymous):

ok that makes a lot more sense i just havent had someone take the time to describe how i actuallly do that process ya know :]

OpenStudy (anonymous):

No problem, glad to help :)

OpenStudy (anonymous):

I say, learn the math and prove the people who say "blonds can't do math" wrong :) Anyway, good luck and glad it was helpful...

OpenStudy (anonymous):

thanks so much hopefully if i need help on other problems i can get someone awesome like you to help me :)

OpenStudy (anonymous):

you can "become a fan" or whatever they call it here... that might help you spot me if I'm online for math help. But hopefully the awesome count here is more than just 1 ;)

OpenStudy (anonymous):

sweet will do